In a recent study on world​ happiness, participants were asked to evaluate their current lives on a scale from 0 to​ 10, where 0 represents the worst possible life and 10 represents the best possible life. The mean response was 5.8 with a standard deviation of 2.2. ​(a) What response represents the 93rd ​percentile? ​(b) What response represents the 55th ​percentile? ​(c) What response represents the first quartile? ​(Round to two decimal places as​ needed.)
Answer #1

a) z score for 93rd percentile = 1.48

Hence,

93rd percentile = 5.8 + 1.48*2.2 = 9.05

b) z score for 55th percentile = 0.13

Hence,

55th percentile = 5.8 + 0.13*2.2 = 6.08

c) z score for first quartile (25th percentile) = -0.67

Hence,

25th percentile = 5.8 - 0.67*2.2 = 4.32